There are 12 red checkers and 3 black checkers in a bag. Checkers are selected one at a time, with replacement. Each time, the color of the checker is recorded. Find the probability of selecting a red checker exactly 7 times in 10 selections. Show your work.
Answer by stanbon(57424)

There are 12 red checkers and 3 black checkers in a bag.
Checkers are selected one at a time, with replacement.
Each time, the color of the checker is recorded.
Find the probability of selecting a red checker exactly 7 times in 10 selections.
-------
Binomial Problem with n = 10 and p = 12/15 = 4/5
-------
P(x = 7) = 10C7(4/5)^7*(1/5)^3 = 0.2013
=================
Cheers,
Stan H.
=================